lim (1+a/x+b/(x)^2)^2x=e^2,  then a and b are?


x->infinity

Grade:Select Grade

1 Answers

Rinkoo Gupta
10 years ago
lim ( 1+a/x+b/x^2)^2x=e^2
x->infinite

e^limx->infinte 2x[1+a/x+b/x^2 -1]=e^2

e^limx->infinite2x[a/x+b/x^2]=e^2

e^limx->infinite2[a+b/x]=e^2
e^2a=e^2
=>2a=2
=>a=1

Thanks & Regards

Rinkoo Gupta

AskIITians Faculty